# DyPE: Dynamic Position Extrapolation for Ultra High Resolution Diffusion

### v1

@@ 4.2 Higher-Resolution Class-to-Image Generation
Table 3: ImageNet results comparing FiTv2 and DyPE (PI, NTK-aware, YaRN) on FiTv2-XL/2. We report FID↓, sFID↓, Inception Score (IS)↑, Precision↑, and Recall↑ at $320^{2}$ and $384^{2}$.
Method FID↓ sFID↓ IS↑ Precision↑ Recall↑
@@
After validating our method on text-to-image generation, we next test whether its consistency gains transfer to the core task of class-conditional generation on ImageNet (Russakovsky et al., 2015). We apply DyPE on FiTv2 (Wang et al., 2024), a flexible DiT trained on multiple resolutions. Specifically, we use the FiTv2-XL/2 variant (675675M parameters), which was trained at a maximum resolution of $256\times 256$, and test it on resolutions $320\times 320$ and $384\times 384$. We consider three baselines—PI, NTK-aware interpolation, and YaRN—against their DyPE-enhanced variants (Dy-PI, Dy-NTK, Dy-YaRN). All models are evaluated on the ImageNet validation set ($50,000$ images). We report FID (Heusel et al., 2017), sFID (Nash et al., 2021), Inception Score (IS) (Salimans et al., 2016), Precision, and Recall (Kynkäänniemi et al., 2019). Quantitative results are reported in Tab. 3, show that, as with FLUX, DyPE consistently improves over all vanilla baselines, with Dy-YaRN achieving the best overall performance. Notably, PI severely underperforms relative to base FiTv2, highlighting its ineffectiveness for image generation due to the loss of high-frequency details.
@@ 5 Related Work
Diffusion Transformers. DiT (Peebles & Xie, 2022) have recently emerged as the leading architecture for diffusion-based text-to-image generation  (Ho et al., 2020; Song et al., 2020). While U-Nets (Ronneberger et al., 2015) underpinned earlier advances (Rombach et al., 2022; Podell et al., 2023; Ramesh et al., 2022), DiTs instead adopt transformer-based backbones that naturally capture global context and scale effectively with model and data size, enabling increasingly capable text-to-image models such as FLUX (Lee et al., 2025), Stable-Diffusion-3 (Esser et al., 2024) and subsequent advances (Gao et al., 2024; Liu et al., 2024a; Chen et al., 2023a; Betker et al., ). Yet, training these architectures on ultra-high resolutions (e.g., 44K and beyond) remains an open challenge due to the quadratic cost of self-attention, which quickly becomes prohibitive in both memory and computation at such resolutions.
@@
Ultra-High Resolution Image Synthesis. Despite this limitation, many works explored fine-tuning diffusion models on higher-resolution  (Liu et al., 2025; Cheng et al., 2025; Hoogeboom et al., 2023; Liu et al., 2024b; Ren et al., 2024; Teng et al., 2023; Zheng et al., 2024; Zhang et al., 2025; Huang et al., 2024), yet these remain limited in their ability to scale to ultra-high resolutions due to the expensive tuning phase. Alternatively, patch-based methods (Bar-Tal et al., 2023; Du et al., 2024; He et al., 2023) aim to reduce costs by stitching generated regions, yet often suffer from duplication and local repetition. Input-level techniques suppress undesired semantics (Lin et al., 2024b; Liu et al., 2024c), but are limited to small artifacts and risk information leakage. More recently, tuning-free methods that synthesize full images without retraining (Qiu et al., 2025; Cao et al., 2024; Haji-Ali et al., 2024; Hwang et al., 2024; Jin et al., 2023; Kim et al., 2025; Lee et al., 2023; Lin et al., 2024a; Zhang et al., 2024) offer a practical alternative, but since all such approaches rely on U-Net backbones, adapting them to DiTs is non-trivial, leaving a critical gap for transformer-based methods capable of true end-to-end ultra-high-resolution generation.
@@
Position Extrapolation Schemes. The challenge of ultra–high-resolution generation in DiTs closely mirrors that of long-context generation in language models, often tackled through advances in positional encoding. RoPE (Su et al., 2021) dominates this space, with extrapolation framed as frequency scaling: PI (Chen et al., 2023b) compresses positions to limit phase drift, while NTK-aware (Peng et al., 2023a) and YaRN (Peng et al., 2023b) rescale frequencies to stabilize low modes and suppress unstable high ones. Inspired by these advances, vision models have begun to adopt these techniques. FiT (Lu et al., 2024) and FiT-v2 (Wang et al., 2024) introduce Vision-PI, Vision-NTK, and Vision-YaRN within DiTs by applying these frequency-scaling techniques independently to the horizontal and vertical axes. While this approach allows for flexible aspect-ratio generation and modest resolution gains, it remains a generic solution that overlooks the low-to-high frequency progression inherent to diffusion. Lumina-Next (Zhuo et al., 2024) incorporates timestep dynamics by interpolating from PI to NTK-aware scaling as denoising advances. Yet, its heavy reliance on interpolation throughout the denoising process suppresses high frequencies, yielding blurry outputs. Our work, instead, directly analyzes the diffusion process frequency progression, leading to a principled approach that preserves fine-grained detail without compromising structural fidelity.
@@ 6 Conclusion
As future work, we aim to pursue even more ambitious resolutions, not only through inference-time scaling but also by incorporating time-dependent positional extrapolation into a light tuning phase. Another avenue is video generation, where these principles benefit both spatial fidelity and temporal coherence through time-aware positional extrapolation.
@@ Appendix B Comparison Between DyPE and Lumina-Next
The frequency allocation strategy behind DyPE is based on two complementary observations made in Sec. 3.1. The first related to the fact that low-frequency modes converge early in the sampling process, whereas the high frequency bands are resolved throughout the process. The second, is related to the trade-off exiting extrapolation method must take when trying to capture the entire spectrum of the larger resolution using the the fixed number of representable modes in the mode, $D$. Thus, rather than pinpointing both ends of the spectrum at all times and accommodating these trade-offs, DyPE, exploits the fact hat low-frequencies are resolved earlier to better represent the higher bands and reduce the extrapolation compression.
@@
The results in Tab. 4 show that Dy-NTK-aware achieves the best performance across all metrics. Additionally, a qualitative comparison provided in Fig. 8
@@ Appendix C Implementation Details
Unless otherwise stated, all experiments are conducted on a single L40S GPU. We set $\alpha=1$, $\beta=32$, and use an effective resolution of $L=1024$. Diffusion inference is performed with 28 sampling steps. For our method, we apply $\lambda_{s}=\lambda_{t}=2$.
@@ Appendix E Additional Results
Additional Baseline Comparison.
For completeness, we conduct an additional comparison, evaluating DyPE, specifically Dy-YaRN, against FreeScale (Qiu et al., 2025). Although FreeScale is built on SDXL, an older model than FLUX (which underlies our method), we include it since it is a recent leading approach for ultra-high-resolution image generation ($4096\times 4096$). We run both methods on the Aesthetics-4K benchmark and report CLIPScore, ImageReward, Aesthetic-Score, and FID. The results are summarized in Tab. 7. As shown, DyPE achieves superior performance across all reported metrics.
Table 7: Comparison of DyPE (Dy-YaRN) with FreeScale for $4096\times 4096$ image generation on the Aesthetics-4K benchmark.
@@
FreeScale 25.91 -1.19 5.75 259.24
Dy-YaRN 29.28 1.09 6.67 186.00

### v2

@@ 4.2 Higher-Resolution Class-to-Image Generation
Table 4: ImageNet results on FiTv2-XL/2 comparing PI, NTK, TASR, YaRN and our DyPE variants. We report FID↓, sFID↓, Inception Score (IS)↑, Precision↑, and Recall↑ at $320^{2}$ and $384^{2}$.
FID↓ sFID↓ IS↑ Precision↑ Recall↑
@@
After validating our method on text-to-image generation, we next test whether its consistency gains transfer to the core task of class-conditional generation on ImageNet (Russakovsky et al., 2015). We apply DyPE on FiTv2 (Wang et al., 2024), a flexible DiT trained on multiple resolutions. Specifically, we use the FiTv2-XL/2 variant (675675M parameters), which was trained at a maximum resolution of $256\times 256$, and test it on resolutions $320\times 320$ and $384\times 384$. We compare the standard extrapolation methods (PI, NTK-aware, YaRN) and TASR against our DyPE variants (Dy-PI, Dy-NTK, Dy-YaRN). All models are evaluated on the ImageNet validation set ($50,000$ images). We report FID (Heusel et al., 2017), sFID (Nash et al., 2021), Inception Score (IS) (Salimans et al., 2016), Precision, and Recall (kynkäänniemi2019improved). Quantitative results are reported in Tab. 4, show that, as with FLUX, DyPE consistently improves over all vanilla baselines, with Dy-YaRN achieving the best overall performance. Notably, PI severely underperforms relative to base FiTv2, highlighting its ineffectiveness for image generation due to the loss of high-frequency details.
@@ 5 Related Work
Diffusion Transformers. DiT (Peebles and Xie, 2022) have recently emerged as the leading architecture for diffusion-based text-to-image generation  (Ho et al., 2020; Song et al., 2020). While U-Nets (Ronneberger et al., 2015) underpinned earlier advances (Rombach et al., 2022; Podell et al., 2023; Ramesh et al., 2022), DiTs instead adopt transformer-based backbones that naturally capture global context and scale effectively with model and data size, enabling increasingly capable text-to-image models such as FLUX (Labs, 2024), Stable-Diffusion-3 (Esser et al., 2024b) and subsequent advances (Gao et al., 2024; Liu et al., 2024a; Chen et al., 2023a). Yet, training these architectures on ultra-high resolutions (e.g., 44K and beyond) remains an open challenge due to the quadratic cost of self-attention, which quickly becomes prohibitive in both memory and computation at such resolutions.
@@
Ultra-High Resolution Image Synthesis. Despite this limitation, many works explored fine-tuning diffusion models on higher-resolution  (Liu et al., 2025; Cheng et al., 2025; Hoogeboom et al., 2023; Liu et al., 2024b; Ren et al., 2024; Teng et al., 2023; Zheng et al., 2024; Zhang et al., 2025a; Huang et al., 2024), yet these remain limited in their ability to scale to ultra-high resolutions due to the expensive tuning phase. Alternatively, patch-based methods (Bar-Tal et al., 2023; Du et al., 2024a; He et al., 2023) aim to reduce costs by stitching generated regions, yet often suffer from duplication and local repetition. Input-level techniques suppress undesired semantics (Lin et al., 2024b; Liu et al., 2024c), but are limited to small artifacts and risk information leakage. Complementary strategies improve high-resolution fidelity within U-Net architectures by modifying internal feature processing, such as FreeU (Si et al., 2024), which enhances skip-connection feature mixing, and FAM-Diffusion (Yang et al., 2025), which introduces frequency modulation for sharper high-resolution outputs. More recently, tuning-free methods that synthesize full images without retraining (Qiu et al., 2025; Cao et al., 2025; Haji-Ali et al., 2024; Hwang et al., 2024; Jin et al., 2023; Kim et al., 2025; Lee et al., 2023; Lin et al., 2024a; Zhang et al., 2024) offer a practical alternative, but since all such approaches rely on U-Net backbones, adapting them to DiTs is non-trivial, leaving a critical gap for transformer-based methods capable of true end-to-end ultra-high-resolution generation.
@@
Position Extrapolation Schemes. The challenge of ultra–high-resolution generation in DiTs closely mirrors that of long-context generation in language models, often tackled through advances in positional encoding. RoPE (Su et al., 2021) dominates this space, with extrapolation framed as frequency scaling: PI (Chen et al., 2023b) compresses positions to limit phase drift, while NTK-aware and YaRN (Peng et al., 2023) rescale frequencies to stabilize low modes and suppress unstable high ones. Inspired by these advances, vision models have begun to adopt these techniques. FiT (Lu et al., 2024) and FiT-v2 (Wang et al., 2024) introduce Vision-PI, Vision-NTK, and Vision-YaRN within DiTs by applying these frequency-scaling techniques independently to the horizontal and vertical axes. While this approach allows for flexible aspect-ratio generation and modest resolution gains, it remains a generic solution that overlooks the low-to-high frequency progression inherent to diffusion. RIFLEx (Zhao et al., 2025) demonstrates that frequency-aware extrapolation can also be effective in DiTs for video, enabling substantial temporal length extension. However, RIFLEx focuses exclusively on the temporal axis and does not address spatial resolution scaling. Lumina-Next (Zhuo et al., 2024) incorporates timestep dynamics by interpolating from PI to NTK-aware scaling as denoising advances. Yet, its heavy reliance on interpolation throughout the denoising process suppresses high frequencies, yielding blurry outputs. Our work, instead, directly analyzes the diffusion process frequency progression, leading to a principled approach that preserves fine-grained detail without compromising structural fidelity.
@@ 6 Conclusion
As future work, we aim to pursue even more ambitious resolutions, not only through inference-time scaling but also by incorporating time-dependent positional extrapolation into a light tuning phase.
Impact Statement
This paper presents work whose goal is to advance the efficiency and scalability of diffusion transformers for ultra-high-resolution image synthesis. Our research does not involve human subjects, personal data, or deployment in user-facing systems. All experiments were conducted on widely adopted, publicly available benchmarks.
We acknowledge that advancements in high-fidelity image generation carry potential societal consequences, particularly regarding the risk of misuse for generating deceptive or harmful content. However, our contributions are strictly focused on methodological improvements and evaluation on standardized datasets. We believe that improving the efficiency of these models does not inherently increase these risks beyond those already well-established in the field of generative modeling. We are committed to the principles of responsible AI research and transparency in methodology.
@@ Appendix B Comparison Between DyPE and Lumina-Next
The frequency allocation strategy behind DyPE is based on two complementary observations made in Sec. 3.1. The first related to the fact that low-frequency modes converge early in the sampling process, whereas the high frequency bands are resolved throughout the process. The second, is related to the trade-off exiting extrapolation method must take when trying to capture the entire spectrum of the larger resolution using the the fixed number of representable modes in the mode, $D$. Thus, rather than pinpointing both ends of the spectrum at all times and accommodating these trade-offs, DyPE, exploits the fact that low-frequencies are resolved earlier to better represent the higher bands and reduce the extrapolation compression.
@@
The results in Tab. 5 show that Dy-NTK-aware achieves the best performance across all metrics. Additionally, a qualitative comparison is provided in Fig. 8
@@ Appendix C Implementation Details
Unless otherwise stated, all experiments are conducted on a single L40S GPU. We set $\alpha=1$, $\beta=32$, and use an effective resolution of $L=1024$. Diffusion inference is performed with 28 sampling steps. For our method, we apply $\lambda_{s}=\lambda_{t}=2$. Code will be released upon acceptance.
@@ Appendix E Additional Results
Evaluation on Additional DiT-Based Models.
To demonstrate the generalization capabilities of our approach beyond FLUX, we implemented DyPE on Qwen-Image (Wu et al., 2025). We compared the vanilla model, the static YaRN baseline, and our dynamic variant, Dy-YaRN. All methods were evaluated on the Aesthetic-4K benchmark using the metrics established in our main experiments. As shown in Tab. 8, Dy-YaRN achieves the best performance across all metrics (CLIPScore, ImageReward, Aesthetic-Score, and FID), validating the robustness and effectiveness of our method across different model architectures. Qualitative comparisons, illustrated in Fig. 9, further confirm that Dy-YaRN produces superior visual quality with fewer artifacts compared to the baselines.
Table 8: Comparison of Qwen-Image, YaRN, and Dy-YaRN on the Aesthetic-4K benchmark.
@@
Qwen-Image (Vanilla) 27.98 -0.26 5.52 201.15
Qwen-Image + YaRN 28.71 0.60 6.17 199.24
Qwen-Image + Dy-YaRN 28.85 0.74 6.20 197.38
Figure 9: Qualitative comparison of Qwen-Image, YaRN, and Dy-YaRN (both adapted on top of Qwen-Image) at a resolution of $4096\times 4096$.
High-Resolution Video Generation.
We extended our evaluation to the video domain by applying DyPE to the Wan2.1 1.3B model (Wan et al., 2025). While the original model has an effective resolution of $832\times 480$, we assessed generation capabilities at a higher resolution of $1280\times 720$. Due to GPU memory constraints, we fixed the sequence length to 33 frames. We compared our approach, specifically Dy-YaRN, against the vanilla Wan2.1 model and YaRN using the VBench benchmark (Huang et al., 2023).
The quantitative results are summarized in Tab. 9. Dy-YaRN outperforms the vanilla model across all categories. Notably, while YaRN suffers a degradation in the Imaging Quality score compared to the vanilla baseline, our dynamic approach improves it. Figure 10 provides a qualitative comparison.
Table 9: Comparison of Wan2.1, YaRN, and Dy-YaRN on high-resolution video generation ($1280\times 720$) using VBench.
Method Subj. Const. $\uparrow$ Bg. Const. $\uparrow$ Mot. Smooth. $\uparrow$ Dyn. Deg. $\uparrow$ Aesth. Qual. $\uparrow$ Img. Qual. $\uparrow$
Wan 2.1 (Vanilla) 0.9170 0.9518 0.9791 0.6532 0.4035 0.5107
Wan 2.1 + YaRN 0.9262 0.9513 0.9868 0.7350 0.4979 0.4364
Wan 2.1 + Dy-YaRN 0.9303 0.9536 0.9899 0.8023 0.5095 0.6127
Figure 10: Qualitative comparison between Wan 2.1, YaRN, and Dy-YaRN, both on top of Wan 2.1, applied to video generation at $1280\times 720$.
High-Resolution Image Editing.
To demonstrate DyPE’s versatility, we further integrated it with the image editing model Qwen-Image-Edit-2509 (Wu et al., 2025). We evaluated performance on multi-concept composition, a challenging task requiring the seamless integration of distinct reference objects into a unified high-resolution scene. Experiments were conducted at an ultra-high resolution of $2656\times 2656$. As shown in Fig. 11, the baseline (vanilla Qwen-Image-Edit-2509) tends to suffer from object duplication, such as the cats and baskets. In contrast, DyPE effectively mitigates these repetition artifacts, ensuring more precise object integration.
Figure 11: Qualitative comparison of high-resolution multi-concept composition at $2656\times 2656$ between Dy-YaRN and the vanilla Qwen-Image-Edit-2509.
@@
Qualitative Comparison with General Baselines.
Building on the comparisons presented in Sec. 4.1.2, we further provide qualitative results that highlight the differences between our approach and existing baselines (see Fig. 14). e
Figure 14: Qualitative results at $4096^{2}$ resolution using two representative prompts from Aesthetic-4K. We compare DemoFusion, Diffusion-4K, HiFlow, Dy-YaRN+HiFlow and Dy-YaRN.
@@
Appendix F Attention Entropy Analysis
Recent work (Jin et al., 2023) suggests that a primary reason trained attention mechanisms fail to generalize to higher resolutions is the shift in attention entropy relative to the training distribution. To investigate this, we analyze the Normalized Attention Entropy (scaled by the logarithm of the sequence length) as a function of the diffusion timestep, averaged across all layers and heads. We conducted this analysis using 20 random prompts from the Aesthetic-4K dataset, comparing the baseline FLUX model at its native resolution ($1024\times 1024$) against FLUX, YaRN, and Dy-YaRN at a resolution of $4096\times 4096$.
As illustrated in Fig. 18, Dy-YaRN best preserves the attention structure of the original distribution. Quantitatively, in terms of deviation from the baseline profile (measured by Mean Absolute Error), Dy-YaRN achieves the lowest deviation ($0.0455$), outperforming both YaRN ($0.0476$) and the vanilla model ($0.0529$). This confirms that DyPE effectively mitigates the entropy shift typically observed during resolution extrapolation.
Figure 18: Deviation of Normalized Attention Entropy from the baseline ($1024\times 1024$) profile across diffusion timesteps. Lower values indicate better preservation of the original attention structure.